# miniCTX: Neural Theorem Proving with

### v1

@@ 3.1 Core challenges: dependencies, length, constraints, and varied content
Helpful context. We also note that context may make some theorems easier to prove compared to standalone theorems such as competition problems. For instance, Lean source code can have natural language comments or Latex proofs that may help constrain the space of possible proofs. Moreover, some proofs within a repository or file often have analogous patterns or structure, which intuitively may make subsequent theorems easier to prove (e.g., see §A.2).
Automatically updating the benchmark. miniCTX’s format is amenable to periodically updating the benchmark with new projects to ensure that theorems, definitions, or full projects are not seen by language models trained prior to a particular date. Future updates will be automatically extracted from new Lean projects and commits using ntp-toolkit (§4.1). Users can also create their own evaluation data with ntp-toolkit. To test a model’s ability to adapt to in-file contexts, we recommend using files that are long, but do not contain intra-repository imports. For example, Rectangle.lean from the Prime Number Theorem contained 47 theorems and only imported Mathlib packages.
4 File-tuning for theorem proving with context
Writing a proof can be seen as a sequential process $(x_{1},y_{1}),(x_{2},y_{2}),\ldots$ of states $x_{t}$ and tactics $y_{t}$. A state contains what is left to prove (the goal), and available information (the hypotheses). A tactic transitions the proof to a new state. If the state contains no remaining goals, the proof is complete. Concretely, a user applies tactics by writing Lean code, Lean keeps track of the state, and the development environment shows the state and the written code.
@@ 4 File-tuning for theorem proving with context
The traditional approach to training a language model for theorem proving is to train a model on (state, tactic) pairs, i.e., train it to predict the next step of a proof (i.e., the tactic), given the state provided by the proof assistant (i.e., the proof state) [9, 7, 19, 8]. A drawback to this approach is that at test time, the model is not aware of new context outside of the proof state, such as new lemmas. We will see later on that models trained with this state-tactic approach fail at context-dependent proving.
As a stronger baseline for context-dependent proving, we present file-tuning, a simple recipe that trains with (preceding file context, proof state, next-tactic) tuples instead of training with (proof state, next-tactic) pairs (Figure 2). This lets the model use new definitions, theorems, or other information that are defined prior to the current tactic at training or at test time. In practice, file-tuning requires extracting contexts and proof states from Lean. We describe our toolkit for doing so next.
4.1 ntp-toolkit for automated data extraction and evaluation
We provide ntp-toolkit to extract training data, evaluation data, and interactive Lean evaluation.
Data extraction.
ntp-toolkit contains a general-purpose data extraction tool that extracts examples from an arbitrary Lean 4 repository and formats them into examples for language-model fine-tuning. The tool is implemented in Lean based on Kim Morrison’s lean-training-data [20]. Specifically, ntp-toolkit takes in a configuration file with one or more Lean repositories specified. Each repository is transformed into next-tactic and full proof examples stored in JSON Lines files. The next-tactic data is suitable for making file-tuning examples of the form (context, state, next-tactic). The full proof data is suitable for making evaluation or training examples of the form (context, theorem, proof). See Appendix §C.1 for the JSON formats provided by the data extraction.
Instruction-tuning data.
ntp-toolkit re-formats extracted data into input-output examples for file-tuning and state-tactic tuning. Each example is preceded with a natural language instruction that describes the task (generating a next-tactic) and the available inputs (e.g., file contents, proof state). Data of this form is often referred to as “instruction-tuning data”. The specific instructions and formats that we used are shown in the Appendix (§C.2). Users can also adjust the formats as needed.
Interaction via Lean REPL.
Proving a theorem with a language model requires interacting with Lean: sending tactics, receiving error messages, etc. In an effort to better integrate LLMs and Lean, we developed a Python wrapper for the Lean Read-Eval-Print Loop (REPL) [17]. The Lean REPL offers an interactive environment for submitting code to and receiving messages from a Lean instance. We provide a Python interface for (1) submitting both complete Lean files and individual tactics, and (2) receiving and handling feedback from Lean, such as error messages or indicators that a tactic application is valid or a proof is complete. This interface enables standard tactic-based proving along with other strategies (e.g., submitting and checking a full proof).
@@ 5 Experiments
First, we describe our procedure for file-tuning a language model on examples extracted from Mathlib. Second, we evaluate the file-tuned model and other baselines on miniCTX. We study performance along various axes, such as dependencies, context length, difficulty, and the type of information in the context. Our experiments show that traditional state-tactic models struggle in the context-dependent setting compared to file-tuned models that can use context at proving time. Finally, our investigation reveals several open challenges and future directions that we discuss in §5.5
5.1 File-tuning on Mathlib
Data extraction.
@@ 5.1 File-tuning on Mathlib
We ran ntp-toolkit’s next-tactic extraction on a snapshot of Mathlib, yielding 307,049 examples available at l3lab/ntp-mathlib. We then ran ntp-toolkit’s instruction tuning script on these examples, yielding file-tuning examples and state-tactic examples. For the file-tuning examples, as an initial method for handling the long Lean files, we either truncate the middle of an input file so that the file contents is 1024 tokens, or take only the preceding 1024 tokens, with the strategy selected at random for each example. The state-tactic examples are at l3lab/ntp-mathlib-instruct-st. The union of file-tuning and state-tactic examples are at l3lab/ntp-mathlib-instruct-context, split into 583k train, 15k dev, and 15k test examples.
File-tuning.
Next, we fine-tune a language model on the union of file-tuning and state-tactic examples. We use the DeepSeek-Coder-1.3b language model [21] based on its performance on code generation tasks and its size, which allowed us to fine-tune with our computational budget. We fine-tune the model for 3 epochs and select the model based on held-out validation perplexity evaluated every 4,000 steps. The model is available at l3lab/ntp-mathlib-context-deepseek-coder-1.3b.
State-tactic tuning. We fine-tune a similar model on l3lab/ntp-mathlib-instruct-st using the same hyperparameters. The model is available at l3lab/ntp-mathlib-st-deepseek-coder-1.3b.
5.2 Evaluation setup
We evaluate models for the task of tactic-based theorem proving: given a theorem statement, a model generates one tactic at a time while receiving states from the proof assistant. We use a standard best-first search strategy [9, 7, 22, 8, 1] which prioritizes partial proofs based on the model’s average log probabilities. This search method is parameterized by the number of generated tactics per iteration $S$, and the maximum number of iterations $T$. We use the setting from [23, 1] ($S=32$, and $T=100$).
We evaluate five types of baselines: (1) pass@1 full proof generation using GPT-4o: we prompt the model with only the theorem statement and require it to generate a complete proof (see Appendix (§C.2) for details of the prompts and few-shot examples); (2) pass@1 full proof generation with in-file context using GPT-4o: we supplement the theorem statement with up to 8000 tokens of in-file context; (3) the file-tuning model described in (§5.1); (4) the state-tactic model described in (§5.1); and (5) a state-tactic prompting model: we prompt a pre-trained language model with (state, tactic) examples, following [23]. We use the Llemma-7b model [23].
5.3 Top-level results
File-tuning maintains performance on competition problems.
First, we evaluate on miniF2F, a standard benchmark based on competition problems that do not require the use of context. We adapt it to the context-dependent setting by simply providing the model with the miniF2F import statements. As seen in Table 2, the state-tactic models achieve 32.79%, which is higher than existing methods tuned on LeanDojo [8] (e.g., 26.50% in  [8], 27.90% in  [1]). This validates the efficacy of ntp-toolkit’s data extraction and our fine-tuning. The file-tuned model achieves 33.61%, showing that file-tuning can maintain or even improve performance in a setting that does not require context.
File-tuning substantially improves context-dependent proving.
Second, Table 2 shows results on miniCTX. These proofs typically depend on more context than those in miniF2F. We see a dramatic improvement for the file-tuned model over the state-tactic methods. The file-tuned model achieves 31.65% compared to 9.31% from the state-tactic model. The gap shows the importance of evaluating the context-dependent setting and theorem proving methods that use context beyond the proof state.
@@ 5.3 Top-level results
Models Test Prime PFR Mathlib HTPI Avg.
GPT-4o (full proof) - 1.15% 5.56% 2.00% 9.73% 5.59%
GPT-4o (+ context) - 13.79% 1.85% 18.00% 31.89% 22.07%
State-tactic prompting 28.28% 19.54% 5.56% 16.00% 19.15% 20.61%
State-tactic tuning 32.79% 11.49% 5.56% 22.00% 5.95% 9.31%
File tuning 33.61% 32.18% 5.56% 34.00% 38.38% 31.65%
Table 2: Performance comparison of different models on MiniF2F and MiniCTX.
Figure 3: Performance partitioned by dependency type. File-tuned models substantially outperform state-tactic tuned models on theorems with definition and/or theorem dependencies.
@@ 5.4 Analysis
We analyze the file-tuned and state-tactic models on miniCTX further along several axes, including the kinds of contextual dependencies, the difficulty, and the content made available in the context.
File-tuning especially helps on problems with dependencies.
@@
We use the miniCTX metadata to categorize theorems from Prime, PFR and Mathlib splits based on their in-file dependencies. Figure 3 shows the performance on problems that depend on in-file definitions, in-file theorems, both, or neither. We also show miniF2F as an additional reference point for problems without in-file dependencies. The file-tuned model shows a marked improvement over the state-tactic tuned model in all scenarios that have dependencies (outlined in black). The gap is much larger than on problems without dependencies. We conclude that file-tuning specifically helps in the realistic setting of proving with new definitions and theorems.
Some definition dependencies can be handled given access only to the proof state.
Interestingly, the state-tactic tuned model performs decently on definition-only problems, achieving a success rate of 27.66% despite lacking direct access to explicit definitions. Upon closer inspection, we recalled that the goal state provided by Lean includes the names and types of definitions, allowing the model to call automation tactics such as simp with correct arguments (See Appendix §D.5).
Context Type Accuracy (%)
No Context 17.02
Imports & Definition 29.79
Theorems 38.29
No proof 46.81
All 46.81
Table 3: Ablating components of the context.
Definitions and theorems in the context are both important.
To determine the relative importance of various context components, in Table 3 we evaluated the file-tuned model’s performance under two modified settings: one where only in-file definitions are provided, and another where only in-file theorems are included. Performance drops when the model is restricted to using either component in isolation compared to the full context, which includes both definitions and theorems.
Suprisingly, pass rate was similar without proofs in the context.
We also investigate the impact of proofs of previous theorems in the context. We modified the theorem proving context of Rectangle.lean file by replacing all proofs associated with theorems in the context with the placeholder “sorry”, while retaining the theorem statements. Surprisingly, the pass rate remained the same as the original setting in which full proofs were available (though, the models have slightly different sets of successful problems) However, we inspected some proofs and found that our model did have some ability to replicate the proof strategies outlined in the context, as illustrated in Appendix §D.3.
File-tuning improves across all difficulty levels and context lengths.
@@
Finally, Appendix §D.2 shows performance on problems categorized by the length of the human-written proof (when available), which we take as a rough proxy of the problem difficulty. The file-tuned model improved on all three difficulty categories. Appendix §D.2 also shows that file-tuning had improved accuracy across context lengths, particularly for problems with longer contexts. Longer contexts may imply more dependencies, suggesting that these problems can benefit more from file-tuning.
Models rely on common symbolic automation.
@@
To demonstrate an additional kind of context-dependence, we perform an additional analysis on Math2001 [15], which contains the Lean code for the book The Mechanics of Proof by Heather Macbeth. [^4] See Appendix B.1 for further details on Math2001. Due to licensing we do not include it in miniCTX. In particular, the textbook code disables powerful automation tactics including simp and nlinarith to promote manual reasoning, akin to traditional textbook exercises. For example, Math 2001 includes numerous arithmetic problems that are trivial with automation tactics (e.g., nlinarith) but are challenging for models to explicitly prove with step-by-step reasoning (e.g., via calc). In Table 5 we evaluate models with the automation disabled, and observe substantial performance drops, confirming a heavy reliance on these automation tactics. An examination of the training corpus further revealed a general dependency on automated tactics within real Lean projects, indicating that our models have learned to rely on these tactics.
5.5 Discussion and future challenges
@@
Making better use of long-contexts. Our file-tuning method simply truncates contexts to be within a token budget (1024 in our experiments), which can discard useful contextual information. We found gains in providing GPT-4 8,000 tokens of context compared to not providing it context, but its absolute performance was still low. There are several possible strategies that can be explored in future work, including feeding in the entire context, retrieval, or mixtures of the two.
Repository-level context. We focused on evaluating in-file context in this paper. As shown in §D.1, many problems require using context outside of the current file. miniCTX is amenable to studying the cross-file context setting as well. It provides annotations for when a theorem uses context outside of the file, and repository metadata that can be used to support cross-file context.
Challenging proofs. Finally, using context through file tuning did not improve performance on the challenging PFR proofs. Moreover, performance is relatively low (19%) on proofs that had a human-written proof of longer than five lines. Proving these kinds of theorems is an open problem.
Working with constraints. As we showed in Table 5, model performance drops when the proof cannot use standard automation tactics. This can have implications in educational settings such as those in Math 2001, where for instance a more explicit step-by-step proof is desired. Improving performance in this setting of miniCTX is an interesting future direction.

### v2

@@ 3.3 Key features and challenges
Beyond previous datasets like LeanDojo (Yang et al., 2023) and CoqGym (Yang & Deng, 2019), which include relevant definitions and theorems, miniCTX includes additional useful contextual information that may make some theorems easier to prove compared to standalone theorems. For instance, Lean source code can have natural language comments that may help constrain the space of possible proofs. Moreover, some proofs within a file often have analogous patterns or structure, which may make subsequent theorems easier to prove (see §A.2). These additional forms of context occur in the real-world process of formalization, yet their use in neural theorem proving is underexplored.
Automatically updating the benchmark. Most modern neural theorem provers use a large language model as a backbone. Therefore, it is crucial to ensure that evaluation content is not seen during (pre-)training, a problem not addressed by previous benchmarks. miniCTX’s format is amenable to periodically updating the benchmark with new projects to ensure that proofs are not seen by language models trained prior to a particular date. Future periodic updates will be automatically extracted from new Lean projects and commits using ntp-toolkit (§3.4). See Figure 2 for an illustration.
Figure 2: miniCTX is automatically updated with Lean projects to stay ahead of LLM training cutoff dates, making it a suitable benchmark for real-world theorem proving for pre-trained models.
@@
3.4 NTP-Toolkit: automated data extraction and evaluation
We provide ntp-toolkit that automatically extracts miniCTX and metadata. ntp-toolkit is also used to extract data for training baselines, and it provides an interactive Lean evaluation environment.
Data extraction. ntp-toolkit contains a general-purpose data extraction tool that extracts examples from an arbitrary Lean 4 repository and formats them into individual theorems in miniCTX. The tool is implemented in Lean based on lean-training-data (Morrison, 2023). Specifically, ntp-toolkit takes in a configuration file with some Lean repositories specified. Then for each theorem in each repository, it outputs a JSON-formatted entry in miniCTX, including the full context information, theorem statement, and proof. It also extracts useful metadata that is used for automated problem selection in miniCTX (see §3.2), and for our experimental analysis. See Appendix §A and §C.1 for the formats provided by the data extraction.
Interaction via Lean REPL. In order to evaluate a model on miniCTX, a model needs to efficiently communicate generated proofs with Lean and receive current state and error information. To better integrate LLMs and Lean, we developed a Python wrapper for the Lean Read-Eval-Print Loop (REPL) (Lean Prover Community, 2024). The Lean REPL offers an interactive environment for submitting code to and receiving messages from a Lean instance. We provide a Python interface for (1) submitting both complete Lean proofs and individual tactics, and (2) receiving and handling feedback from Lean, such as error messages or indicators that a tactic application is valid or a proof is complete. This interface enables simpler evaluation on the miniCTX benchmark.
@@ 4 Experiments
We evaluate several baselines on miniCTX, demonstrating the importance of context in real-world theorem proving. We study performance along various axes, including premise dependency, context length, difficulty, and the type of information in the context. Our experiments show that file-tuned models, which can utilize context at evaluation time, improve drastically over traditional state-tactic models in the context dependent setting. Moreover, we discover that this real-world performance boost cannot be readily measured by existing benchmarks such as miniF2F. Our investigation reveals several open challenges that we discuss in §5.
@@
Training data. We ran ntp-toolkit’s next-tactic extraction on a 2023 snapshot of Mathlib, yielding 307,049 examples. We then ran ntp-toolkit’s instruction tuning script on these examples, yielding training examples and for state-tactic and file-tuning (see §4.1). For the file-tuning examples, as an initial method for handling the long Lean files, we either truncate the middle of an input file so that the file contents is 1024 tokens, or take only the preceding 1024 tokens, with the strategy selected at random for each example. We open-source our training data for both file-tuning and state-tactic tuning, split into 583k train, 15k dev, and 15k test examples.
4.1 Baselines
We select several baselines that we evaluate on miniCTX, as follows:
Prompting LLMs. We first test the ability of a state-of-the-art API-based model, GPT-4o, to generate the complete proof given a theorem statement in Lean, with several few-shot examples provided for guidance. We generate 8 such proof samples and measure pass@8 proof rate. We also test whether adding context in the form of preceding file contents or retrieved premises improves the proof rate.
State-tactic prompting. Another common approach to theorem proving using language models is to let the model generate a tactic given the current proof state (Han et al., 2022; Yang et al., 2023; Polu & Sutskever, 2020; Lample et al., 2022). Therefore, we test the state-tactic prompting setting, which prompts a model specialized for mathematical tasks, Llemma-7b (Azerbayev et al., 2024), to output a tactic given a proof state. At test time, the model generates one tactic at a time, and we use a best-first search to construct full proofs (Han et al., 2022; Yang et al., 2023; Polu & Sutskever, 2020).
State-tactic tuning. We can further improve tactic generation by fine-tuning language models on human-written (state $x_{t}$, tactic $y_{t}$) pairs. We follow this state-tactic framework and fine-tune a state-tactic tuned model from DeepSeek-Coder-1.3b (Guo et al., 2024) to input proof states and output tactics, trained on human-written tactics in Mathlib, the main mathematical library in Lean, extracted by ntp-toolkit. Similarly, we use best-first search at test time.
File-tuning. A drawback to the setups above is that they only consider static (state, tactic) pairs, and do not take into account new context, including comments, definitions, lemmas, and file structure, encountered during test time. Therefore, we test whether supplying context $c$, in the form of preceding file contents, to the model improves performance. Similar to state-tactic tuning, we fine-tune a 1.3b model on human-written (state $x_{t}$, context $c$, tactic $y_{t}$) triples to generate tactics based on the current context and proof state, resulting in the file-tuned model (see Figure 3 for an illustration).
Premise selection. While in-file context offers valuable information to models, it still overlooks external resources like imported modules, which are often crucial in real-world interactive theorem proving. To better simulate a complete context and evaluate on project-level generalization, we apply premise selection to extract relevant premises from imported files within the same repository. Specifically, we use the premise retriever provided by LeanDojo  (Yang et al., 2023) to identify the top 20 most relevant definitions or lemmas from imported modules and append them to the in-file context. Given that the models we are evaluating are trained on Mathlib, and they exhibit a strong ability to apply Mathlib lemmas, we did not include Mathlib in the potential premises to ensure that the models gain as much information as possible from the provided context (both infile and crossfile). All potential premises are automatically extracted using our toolkit, ensuring an efficient and automated process.
Table 3: Performance comparison (%) of different models on miniF2F and miniCTX.
@@ 4.1 Baselines
Method Test Prime PFR PFRcross Mathlib HTPI HEP SciLean Avg.
GPT-4o (full proof) 7.06 1.85 6.98 14.00 13.33 31.15 6.52 11.72
   + context 31.76 5.56 34.88 26.00 17.78 49.18 17.39 27.08
   + context + premise 29.41 7.41 39.53 15.56 44.26 21.74 26.82
State-tactic prompting 28.28 20.00 5.56 0.00 16.00 0.00 31.15 19.57 14.58
State-tactic tuning 32.79 17.65 5.56 0.00 22.00 11.11 52.46 19.57 19.53
File tuning 33.61 40.00 5.56 44.19 34.00 15.56 60.66 45.65 35.94
   + premise 42.35 11.11 16.28 8.89 50.82 32.61 30.21
4.2 Results
Context-dependent methods improve theorem proving. Table 3 shows baseline performances on miniCTX. We see a dramatic improvement for the file-tuned model (trained on full file context) over the state-tactic model (trained only on proof states) (35.94% vs. 19.53%). Similarly, providing the preceding file context, which includes definitions and lemmas, to GPT-4o results in dramatic improvement compared to using just the proof state (27.08% vs. 11.72%). These findings highlight the importance of providing models with rich contextual information beyond the immediate proof state, also demonstrating that miniCTX is able to measure this ability of context-dependent proving.
Premise selection improves performance on high cross-file dependency splits. The results in Table 3 indicate that premise selection has a mixed impact on model performance. For the GPT-4o, premise selection improves performance on high cross-file dependency splits, such as PFR, PFRcross, and SciLean. This suggests that premise selection helps capture the cross-file context, enabling GPT-4o to make better use of cross-file information. However, for the file-tuned model, premise selection does not consistently improve results, and even performs worse on the PFRcross split, which was designed to evaluate the effective use of cross-file premises. This suggests that the retrieved premises differ significantly from the in-file context. Therefore, developing methods that effectively support the integration of cross-file context (e.g., premise selection) alongside in-file context remains an interesting open research direction for improving performance on the miniCTX benchmark.
Evaluation on miniF2F. We evaluate baselines on miniF2F, a standard benchmark based on competition problems that do not require context. The file-tuned model improves very little beyond the state-tactic model (33.61% vs. 32.79%), showing that the dramatic difference in context-dependent proving abilities seen on miniCTX cannot be captured by miniF2F.
Figure 4: Model performance by dependency on premises. For each theorem in miniCTX, we record as metadata whether its human-written proof depends on other definitions or theorems in the same file (“in-file”) or in other files (“cross-file”), and test the performance of baselines on each type.
@@ 4.3 Analysis
We analyze the baseline models on miniCTX further along several axes, including the kinds of contextual dependencies, the difficulty, and the content made available in the context.
@@
File-tuning especially helps on problems with infile dependencies. We use the miniCTX metadata to categorize theorems based on their in-file dependencies. Figure 6 shows the performance of state-tactic tuned model and file-tuned model on problems with in-file dependencies compared to those without. We also show miniF2F as an additional reference point for problems without in-file dependencies. The file-tuned model shows a marked improvement over the state-tactic tuned model, especially in problems that have dependencies on context. We conclude that file-tuning specifically helps in the realistic setting of theorem proving with new definitions and theorems in context.
Premise selection helps but may interfere with in-file context. We use miniCTX metadata to categorize problems based on their cross-file dependencies, evaluating the impact of premise selection across the entire dataset. As shown in Figure 4, GPT-4o benefits significantly from premise selection on problems with high cross-file dependencies, showing improved performance when leveraging relevant premises from imported files. However, we also observe that premise selection can interfere with in-file context, leading to inconsistent results, particularly when the available in-file context is relatively short. This suggests that adding cross-file premises may sometimes disrupt the model’s ability to focus on the in-file information. Further analysis of this interference is included in §D.3. This highlights the need for more sophisticated integration strategies that can balance both in-file and cross-file contexts effectively.
Table 4: Ablation study on different context components for theorem proving.
Environment Definitions Lemma Lemma Natural Language File-tuning GPT-4o
Statement Proof Comments
14.12% 8.24%
25.88% 2.35%
24.71% 9.41%
27.06% 22.35%
32.94% 34.12%
28.24% 23.53%
35.29% 31.76%
Models can learn from previous proofs in the file context. To determine the contribution of different components in the in-file context, we conducted an ablation study on the PFR.ForMathlib.Entropy.Basic file, which contains numerous co-related lemmas and rich natural language comments, making it an ideal candidate to investigate the influence of different context components. In this ablation, we systematically removed specific parts of the in-file context and evaluated the model’s ability to generate proofs under these modified conditions. As shown in Table 4, both the file-tuned model and GPT-4o benefit from the inclusion of previous proofs in the file context. This indicates that models are capable of learning proof strategies from existing proofs in the file and effectively applying them to new problems (see §D.4 for more examples).
Natural language comments contribute in certain settings. Our ablation also explored the effect of natural language comments in the in-file context. Though the impact was not dramatic, comments written in natural language were found to be helpful in certain settings. In scenarios where proofs were excluded from the context, adding comments resulted in slight performance gains for both models. For the file-tuned model, these gains were further amplified when proofs were included alongside comments, demonstrating the value of combining formal context with explanatory natural language. However, for GPT-4o, the presence of comments when proofs were included led to a slight decrease in performance, suggesting that effective context selection may vary depending on the model architecture and underlying training characteristics.
@@
File-tuning improves across all difficulty levels and context lengths. Finally, Appendix §D.2 shows performance on problems categorized by the length of the human-written proof (when available), which we take as a rough proxy of the problem difficulty. The file-tuned model improved on all three difficulty categories. Appendix §D.2 also shows that file-tuning had improved accuracy across context lengths, particularly for problems with longer contexts. Longer contexts may imply more dependencies, suggesting that these problems can benefit more from file-tuning.
@@
Models rely on common symbolic automation. To demonstrate an additional kind of context-dependence, we perform an additional analysis on Math2001 (Macbeth, 2023), which is another Lean textbook setting. [^2] See Appendix B.1 for further details on Math2001. Due to licensing we do not include it in miniCTX. In particular, the textbook code disables powerful automation tactics including simp and linarith to promote manual reasoning, akin to traditional textbook exercises. For example, Math2001 includes numerous arithmetic problems that are trivial with automation tactics (e.g., linarith) but are challenging for models to explicitly prove with step-by-step reasoning (e.g., via calc). In Table 6 we evaluate models with the automation disabled, and observe substantial performance drops, confirming the reliance on automation tactics. We also find that the state-tactic tuned model relies on simp for unseen definitions, making it performing similarly well to the file-tuned model on theorems that only rely on new definitions (§D.6).
5 Discussion and future challenges
@@
Making better use of long-contexts. Our file-tuning method simply truncates contexts to be within a token budget (1024), which can discard useful contextual information. We found gains in providing GPT-4o 8,000 tokens of context compared to not providing it context, but its absolute performance was still low. There are several possible strategies that can be explored in future work, including feeding in the entire context, retrieval, or mixtures of the two.
Repository-level context. We focused on evaluating in-file context in this paper. As shown in §D.1, many problems require using context outside of the current file. Although we incorporated premise selection as a means of leveraging cross-file context, our experiments indicate that it does not consistently improve performance, even on datasets with high cross-file dependencies. This suggests a need to further investigate how to better integrate premise selection with in-file context. miniCTX provides sufficient metadata to reconstruct the entire environment, allowing for comprehensive investigation into premise selection and other potential methods for leveraging cross-file context.
Challenging proofs. Using context through file tuning did not improve performance on the challenging PFR proofs. Moreover, performance is relatively low (19%) on proofs that had a human-written proof of longer than five lines (see §D.2). Proving these kinds of theorems remains an open problem.
Working with constraints. As shown in Table 6, model performance drops when the proof cannot use powerful automation tactics. Models have a tendency to invoke these powerful tactics, and struggle with more explicit step-by-step proofs. Improving performance in this setting of miniCTX is an interesting future direction.